OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 1) (5, 4)

OpenStudy (zale101):

(1-4/6-5) = the slope

OpenStudy (zale101):

after you find the slope, plug in to this formula given to you (y = mx + b) m is the slope take one of the two points and plug in x and y y = mx + b

OpenStudy (zale101):

Solving for b will be then easier if you alread have the equation for y, m, and x

OpenStudy (zale101):

btw, i used (y1-y2/x1-x2) to get the slope, which i pluged in the x and ys you provided (1-4/6-5)